a painting job can be completed by 6 painters in 45 days if 24 more painters join the team 10 days after starting work,then how many more days will it take to complete the job
Answer by ptaylor(2198)

Let x=number of additional days it will take after the 24 more painters join in.
If 6 painters can complete the job in 45 days, then the job takes 6*45=270 painter-days to complete. In other one painter can complete the job in 270 days
In 10 days, 6 painters have completed 60 painter-days of work leaving 210 painter-days to be completed
OK now after 10 days, we have 24+6=30 painters to do 210 painter-days of work
So, our equation to solve is:
30x=210
x=7 days---------------additional days it will take after 24 more painters join in
CK
In 7 day, 6 painters do 6*7=42 painter-days of work
In 7 days,24 painters do 7*24=168 painter-days of work
42+168=210
210=210
Hope this helps----ptaylor